Coulomb's Law describes the force between two charged objects. Suppose Object 1 has a charge of +1.50 microcoulombs and Object 2 has a charge of +14.2 coulombs. Object 2 is located a distance of 0.05810 m to the right of object 1. What is the magnitude and direction of the electrical force on object 2 due to Object 1?
Answer #1

Charge on Object 1 = Q = + 1.5 uC = +1.5 * 10^(-6) C
Charge on object 2 = q = +14.2 uC = +14.2 * 10^(-6) C

Distance between the two = r = 0.05810 m

We know, like charges repel each other, so object 1 will push object 2 to further rightwards by force F.

F = kQq/r^(2)

k = 9*10^(9) N.m^(2)/ C^(2)


so, F =[ 9*10^(9) * 1.5 * 10^(-6) * 14.2 * 10^(-6) ]/ (0.05810)^(2)

= 56.79 N rightwards.
Answer #2

since they are the same charge they will repel

so 2 will be pushed to the right

F = k q1 q2/r^2 = 9.0E9*1.5E-6*14.2/0.0581^2=5.68E7 N
Answer #3

electric force = k q1 q2 /r^2

q1 =1.5 * 10^-6 coulumb

q2 = 14.2 coulumb

r=0.0581m

force = k 1.5 * 10^-6 * 14.2/0.0581^2

put the value and get answer

direction is towards right